Grade 11Mechanics

find the dimensions of a and b if f = a root x +bt square .
where f is force x is distance t is time

smriti
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
f = a√x + bt2, where [f] = [M L T −2], [x] = [L] and [t] = [T].
LHS is force. So both the terms on the RHS have the dimensions of force.
[f] = [a√x] = [bt2]
[f] = [a√x]
[M L T −2] = [a L½]
⇒ [a] = [M L½ T −2]
[f] = [bt2]
[M L T −2] = [b T2]
⇒ [b] = [M L T−4]